Consider the case where the predetermined average number of bus arrivals at the Cubao Station is 6 per 15-minute period. Find the pro of at least 4 buses arriving in a 15-minute period.


1
Expert's answer
2021-05-03T04:51:34-0400

Let "X=" the number of buses bus arrivals at the Cubao Station: "X\sim Po(\lambda)."

Given "\lambda=6." Then


"P(X\geq4)=1-P(X<4)"

"=1-P(X=0)-P(X=1)-P(X=2)"

"-P(X=3)=1-\dfrac{e^{-6}\cdot6^0}{0!}-\dfrac{e^{-6}\cdot6^1}{1!}"

"-\dfrac{e^{-6}\cdot6^2}{2!}-\dfrac{e^{-6}\cdot6^3}{3!}\approx0.8488"

The probability of at least 4 buses arriving in a 15-minute period is 0.8488.